OpenStudy (anonymous):

Last two on my study guide that I can't figure out: 1.) Write the equation of the line that satisfies the given conditions. Express the final equation in standard form. Contains the point (-5, 7) and is parallel to the line x − 8y = 6 2.) Write the equation of the line that satisfies the given conditions. Express the final equation in standard form. Contains the point (-4, 2) and is perpendicular to the line 5x − y = 4

OpenStudy (mokeira):

Question 1 \[y=\frac{ x }{ 8 }-\frac{ 3 }{ 4 }\] \[m=\frac{ 1 }{ 8 }\] \[m=\frac{\Delta y }{\Delta x }\] \[\frac{ 1 }{ 8 }=\frac{ y-7 }{ x--5 }\] \[\frac{ 1 }{ 8 }=\frac{ y-7 }{x+5 }\] cross multiply \[8(y-7)=x+5\] \[8y-56=x+5\] \[y=\frac{ x }{8 }+\frac{ 61 }{ 8 }\]

OpenStudy (anonymous):

I submitted that and it said it was wrong. It shows that x-8y=-61 is the answer

OpenStudy (anonymous):

@mokeira

OpenStudy (mokeira):

hmmm....i dont know

OpenStudy (mokeira):

ooh i know...you multiply m by -1